2. In 30 s, the crankshaft of a truck engine operating at 2400 rpm turns through how many radians?
1
Expert's answer
2020-07-10T10:22:26-0400

As far, as the truck engine operating at 2400 rpm performns 2400 revolutions per minute, then it will perform "2400\/2 = 1200" revolutions during 30 s.

One revolution correspondes to "2\pi" radians. Thus, during 30 s the truck engine turns through:


"N = 1200\cdot 2\pi = 2400\pi \approx 7539.8 \space (radians)"

Answer. 2400pi or 7539.8 radians.